### 03 Jan 2024 v2

# Balancing Specialization and Adaptation in a Transforming Scientific Landscape

Abstract
How scientists navigate between the need to capitalize on their prior knowledge by specializing, and the urge to adapt to evolving research opportunities? Drawing from diverse perspectives on adaptation, in particular from institutional change and cultural evolution, this paper proposes a Bayesian model of the evolution of scientists’ research portfolios in response to transformations in their field. The model relies on scientific abstracts and authorship data to evaluate the influence of intellectual, social, and institutional resources on scientists’ trajectories within a cohort of 2 19521952\,1952 195 high-energy physicists between 2000 and 2019. The reallocation of research efforts is shown to be structured by learning costs, thus enhancing the utility of the scientific capital disseminated among scientists. Two dimensions of social capital, namely “diversity” and “power”, have opposite effects on the magnitude of change in scientists’ research interests: while “diversity” disrupts and expands research interests, “power” stabilizes physicists’ research agendas – as does institutional stability. Social capital plays a more crucial role in shifts between cognitively distant research areas.
*lucas.gautheron@gmail.com
ORCID: $0$-0002-3776-3373
1Interdisciplinary Centre for Science and Technology Studies (IZWT), University of Wuppertal, Germany
2Département d’Études Cognitives, École Normale Supérieure, Paris, France
Keywords
adaptation; specialization; cultural evolution; science of science; human mobility; optimal transport
1 Introduction
Scientists are subject to conflicting incentives. On the one hand, they must work within the realm of their expertise, where they can most effectively exploit their prior knowledge and compete with peers; this conservative preference for familiar research topics is at the root of specialization. On the other hand, scientists are simultaneously compelled to revise their research agenda to engage with more promising research areas in order to benefit from more exposure, to secure funding, or to seize job opportunities. Thus, in some instances, specialization is at odds with the need to adapt to the decline of certain research opportunities and the growth of new ones. How do scientists navigate the trade-off between specialization (i.e. the concentration of their intellectual resources within a narrow cognitive range) and adaptation (i.e. the need to adjust these resources to new realities)? This conflict differs from the “essential tension” between “tradition” and “innovation” proposed by [1, ], or that between “exploration” and “exploitation” [2], which have both been explored quantitatively in previous works [3, 4, 5, 6, 7, 8, 9]. First, “adaptation” is not tantamount to innovation or disruption, for it can be a conformist move (e.g. as a result of a bandwagon effect, [10]). Moreover, unlike “exploration”, adaptation is not a search strategy or a method for acquiring information, but rather the convergence towards a new state more congruent with current realities. Disruptions due to breakthroughs in Machine Learning (e.g. generative Large Language Models) or challenges due to climate change make it particularly pressing to understand the factors that enhance scientists’ adaptability to changing circumstances. Therefore, the present paper investigates scientists’ responses to changes in their field (whether it is driven by epistemic or institutional factors), and the effect of their capital (intellectual, social, or institutional) on their ability to adapt. Drawing insights on adaptation from cultural evolution and institutional change, we develop an unsupervised Bayesian approach to analyze changes in scientists’ research agenda while measuring the effect of “capital” (intellectual or social) on their individual trajectories. The model is applied to a cohort of high-energy physicists between the years 2000 and 2020, a time during which the historical driver of progress in the field – particle accelerators – have been contested by emerging astrophysical experiments, thus transforming the landscape of opportunities.
Our approach reveals trends in the field: the boom of dark matter research – fueled by migrations away from the physics of neutrinos and the electroweak sector – and the partial disintegration of string theory into the study of black holes and holographic dualities. More importantly, this analysis also shows that changes in scientists’ research portfolios are structured by learning costs, as scientific communities adapting to new circumstances address an Optimal Transport problem by reallocating research efforts efficiently. Moreover, two dimensions of social capital, namely “diversity” and “power” [11], have opposite effects on change. While “diversity” of social capital (i.e. the extent to which scientists have access to diverse cognitive resources via their collaborators) promotes change and further diversification of scientists’ research interests, “power” (the size of their network) stabilizes their research interests. Social capital has a stronger effect on transfers between research areas that are more cognitively “distant”. Furthermore, institutional stability is associated with more conservative research interests. Overall, we contribute: i) a theoretical account of the features of change in scientists’ research interests; ii) a novel methodological approach that introduces a model of scientists’ trajectories and measures of intellectual capital, social capital, diversity, and power; and finally, iii) empirical evidence from High-Energy Physics (HEP). More generally, this paper addresses the relative lack of empirical works within the body of literature that investigates science as a cultural evolutionary system [12].
In what follows, Sections 1.1 and 1.2 summarizes previous research and lays out the conceptual background on which the analysis rests, and Section 1.3 introduces the context of high-energy physics to which the model is applied. Section 2 elaborates the methodology: the data (2.1), the topic model approach for measuring authors’ research portfolios (2.2), the proposed measures of intellectual and social “capital” (2.3), and the model of scientists’ trajectories (2.4). Section 3 presents the results: i) the transfers of attention from one research area to another due to changes in the scientific landscape; ii) the structuring role of learning costs in the observed patterns of adaptation (3.2), and iii) the effect of capital (intellectual and social) and position stability on physicists’ strategies (3.3).
1.1 Empirical background
Changes in scientists’ research interests have been studied in previous works. For instance, by mapping the trajectories of 103,246 physicists over 26 years using the American Physical Society (APS) dataset and its topic classification (the Physics and Astronomy Classification Scheme® (PACS)), [5, ] demonstrated that a majority of physicists gradually migrate to entirely different topics by the end of their careers while often staying within the same general area. They reveal differences between subfields of physics, such that “exploitation” (i.e. specialization, as opposed to the “exploration” of new topics) is especially prevalent in particle and nuclear physics. Using the same data, [4, ] instead find an exponentially decaying distribution of changes in scientists’ interests. According to them, the observed distribution of changes in research interests can be qualitatively reproduced by a model in which scientists draw combinations of ideas from a gradually drifting pool of topics. Previous works generally emphasize the graduality of change in research topics, in agreement with an earlier observation by [13, ]. Recognizing that scientists typically investigate several research questions in parallel, [13] proposed four mechanisms of gradual change, including “accretion” (a problem is added to their “problem set”), “selective substitution” (one problem is replaced by another), and “selective disengagement” (one problem is neglected). Recently, [6, ] have analyzed the frequency of topic switching throughout scientific careers, and [8, ] found two types of scientists (“exploiters” versus “explorers”) by applying a gravity model to arXiv data.
While these works document the structure of changes in scientists’ interests, they do not relate these transformations to changes in epistemic and institutional context, to scientists’ incentives, or to the resources that scientists can leverage in order to broaden their research opportunities. Particularly, they do not explain differences in the behaviors of physicists, except in terms of their subfields of origin or their academic ages [5, 6]. [7, ] have taken a significant step in this direction. Using the APS dataset, they show that physicists are more likely to explore areas closer to their core specialization or areas to which they are connected via their collaborators. Their work highlights the crucial importance of collaborations in the expansion of research portfolios, especially for the exploration of cognitively distant research areas. However, their work does not address the transformations of scientists research portfolios throughout time – they do not quantify “change” and its magnitude –, and they recognize the need for further longitudinal analyses.
The present paper complements previous works in several ways. First, since the focus is on adaptation strategies, the core of our approach is both comparative (as in [7], and unlike other previous works) and longitudinal (unlike [7], although their paper includes longitudinal robustness checks). Second, this contribution evaluates previously unexplored aspects, such as the choice between expansion or consolidation of research portfolios and the effect of institutional stability on change. Third, this work relates the findings to the epistemic context of the field and its transformations by performing the analysis at a smaller scale (high-energy physics alone, rather than physics as a whole). Fourth, this work does not rely on the APS dataset and PACS categories, on which most previous works depend [4, 5, 14, 7, 9]. A different dataset is employed, and research areas are clustered using a topic model, so that the contributed methodology can in principle be applied to any body of literature (consequently, some of our results are reproduced on a corpus of Computational Linguistics research [15]). Finally, the proposed approach is grounded in theory, by translating concepts such as capital [16, 17] into quantitative constructs, and by exploiting theoretical insights from diverse approaches to “adaptation”.
1.2 Conceptual framework
Scientists’ opportunities are constrained by their prior knowledge, since entering new research areas involves learning costs. Yet, the exploitation of prior knowledge may require adjustments to new realities. Through a gradual process of adaptation, scientists can strategically retain the benefits of “problem retention” (e.g., the exploitation of “accumulated skills and resources” in one area, or “of an established research network”, [13, p. 106]) while progressively investing resources in new research directions. This is illustrated in Figure 1, which represents the research agendas of one scientist during two distinct time periods. Cells indicate the resources exploited by the scientist (e.g., concepts, models, methods, or techniques) and colors indicate to what problem areas this knowledge applied. A resource can be applied simultaneously to different ends, as expressed by the hatched cell; however, resources are often tailored for specific research purposes. The figure also shows how scientists can enter new research areas by repurposing their knowledge to new ends [18, 19]. We call this strategy “conversion”, in reference to the typology of incremental institutional change proposed by [20, ]. Indeed, as shown in previous works on the transformations of high-energy physics facilities to photon science instruments [21, 22, 23], historical institutionalism can account for gradual adaptations with large cumulative effects taking place in response to scientific and technological change [24]. In this paper, we apply the typology of change to individuals rather than organizations.
Not all knowledge can be successfully applied to new research areas, and scientists are more likely to shift resources and attention to domains that are less “cognitively distant” from their previous expertise. Additionally, as illustrated in Figure 1, entering new research areas typically requires “layering” (again, borrowing the terminology from historical institutionalism), that is, the introduction of new concepts, models, or methods, on top of prior knowledge. However, the acquisition of knowledge entails learning costs, although those can be partially avoided by collaborating with experts in the target domain [7]. Another mode of change is displacement, when the replacement of one research area for another involves significant neglect of prior knowledge. This may not be the preferred strategy, since it fails to take advantage of accumulated knowledge. However, certain knowledge may not apply to a new context, or sometimes there might be reasons to suspend a line of research in order to focus on more promising topics.
Resource 1Resource 2Resource 3Resource 4Resource 5Resource 6Resource 7Resource 8ConversionDisplacement
LayeringResource 1Resource 2Resource 3Resource 4Resource 5Resource XResource YResource Z
Figure 1: Schematic illustration of the changes in a scientist’s “research portfolio” over time. Each color represents a different research area. Resources entail any intellectual or methodological assets that a scientist uses to address scientific problems from each research area. “Conversion” repurposes knowledge to new goals. Displacement is the replacement of certain research interests with little or no transfer of prior knowledge. Layering is the introduction of new research interests via the addition of new knowledge.
Scientists manage two kinds of assets when navigating the trade-off between specialization and adaptation: their own prior expertise, and the expertise to which they have access through their social network. We propose to consider both of these as “capital” [17], which designates the different assets that individuals (e.g. scientists) can accumulate and leverage in the competitive context of their field. “Capital” (whether “economic”, “cultural”, “social” or even “symbolic”, cf. [17]) defines the scope of scientists’ opportunities and therefore their ability to revise their research agenda. This paper considers the cultural capital possessed by scientists in the form of scientific knowledge (which will be referred to as “intellectual capital”), and social capital. Measures that represent these concepts will be proposed, and their effect on the magnitude of transfers of attention across research areas will be evaluated. Emphasis will be put on the divide underlined in [11] between two dimensions of social capital, namely “power” (in fact, network size in the present paper) and “diversity” (of cognitive resources). Group diversity is generally recognized as a factor of adaptation in an evolving environment [25, 26, 27] or in the context of problem solving [28]. “Power” is also plausibly associated with higher abilities.
While capital defines scientists’ opportunities, it is not sufficient to explain why scientists do turn to new research areas or not, which also requires understanding actors’ incentives and why they must respond to these incentives. Consequently, the present paper also considers the effect of institutional stability on migrations between research areas, since the need to respond to changes in the epistemic and institutional environment is presumably different for physicists applying to new jobs than it is for tenured professors and researchers.
Before introducing the methodology, the next section elaborates on the relevance of high-energy physics as an example for investigating adaptation in a transforming scientific landscape.
1.3 The case of high-energy physics: navigating a changing epistemic landscape
The field of particle physics has relied on the input of increasingly large particle colliders to achieve progress. These efforts culminated in 2010 with the start of the Large Hadron Collider (LHC), the largest accelerator ever. However, the LHC has found no evidence for anything that was not already predicted by the Standard Model of particle physics, and it is increasingly plausible that no future accelerator could ever find any evidence for new particles, leading to a situation of “crisis” [29]. Although the LHC will continue to take data for years and plans for future accelerators are being discussed [30], some have even speculated that particle physics as we know it has come to an end [31, 32], “the proscenium [being] captured by astrophysics and cosmology” instead [32]. In a recent assessment of the field, physicist Mikhail Shifman claimed that the “pause in accelerator programs we are witnessing now is not necessarily […] the end of explorations at [high energies]”; instead, such explorations “will continue, perhaps in a new form, with novel devices” [33]. Indeed, new experimental opportunities have emerged in parallel. Gravitational Waves astronomy began in 2015, with the first “direct” detection of such waves from a black-hole merger by LIGO [34]. Other opportunities include direct searches of dark matter of astrophysical origin in underground facilities (e.g. LUX and Xenon experiments), and more precise data on the early universe from the Planck observatory. In fact, astrophysics seems to be increasingly replacing particle colliders in citations across experimental and theoretical high-energy physics. [35]. These transformations have unfolded gradually: the LHC started in 2010 but it operated much below its target energy until 2015; Planck released data on the cosmic microwave background between 2010 and 2018; etc. (see Figure 9, Appendix A.1). The present paper investigates how high-energy physicists have adapted in reaction to these changes in the scientific landscape.
2 Methods
2.1 Data
Our source is the Inspire HEP database [36]. It aggregates high-energy physics literature from various sources, including the main scientific publishers and arXiv, and has been used in a number of works [37, 38, 39, 40, 35]. For the literature on HEP, it is more comprehensive than the APS dataset that has been the focus of previous work on the evolution of research interests, because the latter is limited to a few journals [^1] https://journals.aps.org/datasets. The database is free-access and includes semantic data for each paper (mainly their title and abstract, sometimes also keywords and PACS categories), curated authorship and institutional affiliations, lists of references and associated experiments. This database implements both automatized and manual measures for the disambiguation of author names [^2] Besides the use of “advance algorithms” of author-disambiguation, Inspire invites scientists to correct their own publication record on the website (cf. https://twiki.cern.ch/twiki/pub/Inspire/WebHome/INSPIRE_background.pdf, June 2014) , which has previously allowed [39, ] to analyze gender gaps in physicists’ research careers based on these data. Nevertheless, occasional misidentifications remain possible.
The present analysis includes all papers from the categories “Theory-HEP” and “Phenomenology-HEP” (according to the classification of the database which is based on that of arXiv), to which most HEP publications belong. The minority of purely experimental high-energy physics publications are excluded: such papers are typically authored by thousands of collaborators, and authorship data provide no information about individual experimentalists’ specialization. Therefore, this paper documents how theorists and phenomenologists have adapted to the changes outline above. In total, the corpus includes $D=186,162$ articles published between 2000 and 2019.
[…]
2.2 Measuring research agendas
Our strategy for measuring research portfolios relies neither on PACS categories (in contrast to [4, 14, 5, 7]) nor on citation data (unlike [6]). Instead, research areas are extracted with a topic model that recovers latent “topics” within the corpus and their vocabulary distributions [42, 43]. This makes the approach applicable to a wider range of situations for which only textual data (even short texts) are available. Arguably, texts also provide a more direct access to the structure of knowledge. A coarse-grained classification of the literature is performed into $K=20$ broad “topics”. The number of topics is always somewhat arbitrary, but this is not as much an issue of topic models themselves: as [13, ] puts it, “in such analyses [of problem change], empirical findings will in part reflect the defined scope of problem areas”, which is itself arbitrary. In fact, that topic models give some control over the cognitive scale is a feature, rather than a bug [^4] Previous works based on the PACS categories have leveraged the different levels of this hierarchical classification system to investigate different scales..
We use the Embedding Topic Model developed by [44], which relies on pretrained embeddings representations of the keywords thus providing more reliable classifications of heavy-tailed vocabulary distributions. We train the model the abstracts of the $D=186,162$ documents published between 2000 and 2019. Keywords are extracted from the papers’ titles and abstracts by filtering n-grams between one and three words matching certain syntactic expressions that we can expect to carry scientific information (by designating concepts, models, methods etc.), following the procedure described in [35] and [45]. Keywords’ embeddings are learned using a skip-grap model in a $L=50$ dimension space (few dimensions are needed given the small size of the vocabulary, $V=4,751$). We obtain the topics listed in Appendix A.3, Table 1. This table shows their most frequent keywords, as well as the PACS categories that are most correlated with them. Four of the 20 resulting topics regroup keywords that do not clearly refer to any specific research area (e.g. “paper”, “approach”) and correlate poorly with the PACS categories, which suggests their lack of scientific dimension. Consequently, the present analysis only considers the $K=16$ remaining topics that designate actual research areas (as confirmed by their strong tendency to cite themselves, cf. Appendix A.3.3, and their correlation with the PACS classification, cf. Appendix A.3.4).
Then, we derive $n_{dk}$, the amount of keywords in the abstract of $d$ that refer to a research area $k$ [^5] See Appendix A.3.1 for further details on the counting procedure., and consequently $X_{a,k}$, the amount of times keywords (“resources”, i.e. concepts, models, methods, etc.) in relation to research area $k$ have occurred in papers (co-)authored by $a$ in the initial time-period (2000 to 2009). Mathematically, $X_{a,k}=\sum_{d\in P,a\in A_{d}}n_{d,k}$, where $A_{d}$ is the set of authors of a publication $d$, and $P$ is the set of papers published between 2000 and 2009. The matrix $Y_{a,k}$ is derived identically, using publications from the later time period (2015 to 2019). Research portfolios are represented by the normalized distributions $x_{ak}\equiv X_{ak}/\sum_{k^{\prime}}X_{ak^{\prime}}$ and $y_{ak}\equiv Y_{ak}/\sum_{k^{\prime}}Y_{ak^{\prime}}$. They encode how scientists divided their attention between research areas during each period.
2.3 Measuring capital
Evaluting the effect of intellectual and social capital on scientists’ trajectories requires quantitative measures of these concepts. An extensive review of prior usage of Bourdieu’s notion of capital in bibliometrics has been provided by [46, p. 198-200 ], with most references addressing symbolic and social capital, and none addressing cultural capital. In the following, a consistent approach to intellectual and social capital is proposed, from which measures of diversity and power are derived.
2.3.1 Intellectual capital
Intellectual capital is represented by a vector $\bm{I_{a}}=(I_{ak})$ that measures the concentration of $a$’s intellectual resources in each domain $k\in\{1,\dots,K\}$. It is constructed in a way similar to $\bm{x_{a}}$, summing the contribution of keywords dedicated to each research area in the publications of each author between 2000 and 2009, except that publications are now weighed differently depending on their amount of authors. Indeed, publications with fewer co-authors convey more information about each author’s own expertise. The weight is $1\over|A_{d}|$, where $|A_{d}|$ is the amount of authors of publication $d$ [^6] A justification for this weight is that the probability that a given author has been responsible for introducing any particular concept or method present in the paper is $\mathcal{O}(1/|A_{d}|)$. :
$I_{ak}\propto\sum_{d\in P,a\in A_{d}}\dfrac{n_{d,k}}{|A_{d}|}$ (1)
$\bm{I_{a}}$ is normalized, such that $\sum_{k}I_{ak}=1$; therefore, $\bm{I_{a}}$ only captures the ways scientists divide their cognitive resources between research areas, rather than the “absolute magnitude” of their knowledge of each area.
2.3.2 Social capital
Many measures of scientists’ social capital have been proposed [11, 46], the simplest being the amount of collaborators of a scientist (i.e. degree centrality in the co-authorship network). Other measures are related to betweeness centrality, which concerns the extent to which an actor “bridges” a network (e.g. “brokerage”, i.e., the ability of an individual to overcome “structural holes” in a social network, [47]). The difference between these measures matter; for instance, redundant contacts (and sheer network size) can detrimental in some circumstances. [11, ] distinguishes two general approaches to social capital, depending on whether the emphasis is placed on “power” versus “diversity”. They propose alternative measures based on citations that correlate better with scientists’ performance than measures of degree centrality. All the measures discussed in [11] represent social capital by a single scalar; however, a single number misses multiple dimensions of social capital. In fact, according to [16, ], “the volume of social capital possessed by a particular agent […] depends on the extent of the network of links that he can effectively mobilize, and on the volume of capital (economic, cultural or symbolic) possessed by each of those to whom he is linked”. In that respect, social capital can come in different forms depending on the resources being leveraged via one’s network. In the following, we focus on the intellectual dimension of social capital, which we represent by a vector $\bf{S}_{a}$ that encodes how the amount of expertise in each research area is carried is available to scientists via their network. More specifically, we define $\bf{S}_{a}$ as the sum of the intellectual capital of $a$’s collaborators, weighted by the strength of their relationship:
$\bm{S}_{a}\equiv\sum_{c\in\text{co-authors}(a)}w_{ac}\bm{I}_{c\backslash a}$ (2)
$\bm{I}_{c\backslash a}$ is the intellectual capital of $c$, evaluated by excluding papers co-authored with $a$ (in order to disentangle the effect of an author’s own intellectual capital and intellectual capital available to them in social form). Collaborators outside the cohort are taken into account. The weight $w_{ac}$, which represents the strength of the relationship between $a$ and $c$, is defined as:
$w_{ac}\equiv\max_{d|\{a,c\}\subset A_{d}}\frac{1}{|A_{d}|-1}$ (3)
Where $A_{d}$ is the set of the co-authors of a paper $d$. This weighing scheme – inspired from [48] – captures the fact that a paper with, say, two co-authors, signals a stronger relationship between the authors than a publication with a dozen authors. [^7] Assuming that in a collaboration, each author interacts with a constant amount of co-authors in practice (regardless of the total amount of co-authors), then the probability that they had interactions with one specific co-author in particular is $\propto\frac{1}{|A_{d}|-1}$. However, Equation (3) does not take into account the recency and frequency of collaborations.
2.3.3 Diversity and power
Measures of “diversity” (and “power”) can be readily derived from $\bm{I_{a}}$ (and $\bm{S_{a}}$). A common measure of diversity is the Shannon entropy $H$ [^8] The Shannon entropy of the discrete uniform distribution on $\{1,\dots,n\}$ is $\log{n}$, such as that the exponential of the entropy is roughly the effective amount of outcomes for the distribution ($n$).. Let $D(\bm{I_{a}})=\exp{H(\bm{I_{a}})}$ be the diversity of intellectual capital (and $D(\bm{S_{a}})=\exp{H(\bm{S_{a}})}$ that of social capital). Roughly, these are measures of how many research areas scientists have divided their cognitive or social resources among. In the cohort, individuals typically have cognitive resources in several research areas ($\mu_{D(\bm{I_{a}})}=6.1,\sigma=2.2$), and social capital is even more diverse ($\mu_{D(\bm{S_{a}})}=8.5,\sigma=2.2$). In fact, intuitively, scientific collaborations enable individuals to take advantage of their group’s cognitive diversity. Furthermore, $D(\bm{I_{a}})$ and $D(\bm{S_{a}})$ are highly correlated ($R=0.75$); indeed, individuals with more diverse interests are more able to engage with diverse collaborators. Since the diversity of social capital is only expected to enhance individuals’ abilities when it exceeds that of their own intellectual resources, from now on, only excess social capital diversity $D^{\ast}(\bm{S_{a}})$ (defined as the residuals of the linear regression of $D(\bm{S_{a}})$ against $D(\bm{I_{a}})$, by ordinary least squares) is considered999[7, ] noted the difficulty of determining the direction of the causal relationship between social resources and research interests. This approach aims to overcome the very same challenge..
The “power” dimension of social capital is evaluated as the magnitude of social capital:
$P(\bm{S_{a}})\equiv\sum_{k}S_{ak}=\sum_{c\in\text{co-authors}(a)}w_{ac}$ (4)
The magnitude of social capital is therefore the amount of collaborators weighed by the strength of each relationship. Our measures of diversity and power therefore depart from [11] in which diversity is conflated with network size and power with performance. By combining semantic and authorship data, our approach assesses diversity more directly.
For each of these (diversity and power), an alternative measure is also considered for robustness assessment (based on Stirling’s index for diversity and brokerage for power). All these measures are compared in Appendix A.4.
2.4 Modelling trajectories
Physicists’ trajectories are modeled with a hierarchical multinomial logistic model. The intuition underlying the model is illustrated in Figures 1(a),1(b),1(c), and a more formal representation of the model is given in Figure 1(d). The occurrences of keywords that belong to each research are $k$ in papers by $a$ in the late time period, $\bm{Y_{a}}\in\mathbb{R}^{K}$, are assumed to be a function of the initial research portfolio ($\bm{X_{a}}$), thus capturing the intuition illustrated in Figure 1 that research agendas are transformed via strategic transfers of knowledge from one research area to another. $\bm{Y_{a}}$ is also a function of a mixing matrix $\theta_{a}$ that measures the extent to which attention has been redistributed to different research areas. $\theta_{a}$ is drawn from a hierarchical process, thus capturing the “average” behavior of the cohort. It is also a function of covariates ($\bm{Z_{a}}$ in Figure 1(d)), in this case intellectual and social capital, which are hypothesized to affect how individuals have revised their research agenda.
Formally speaking, we assume that $\bm{Y_{a}}$ derives from a multinomial process involving linear combinations of $(x_{ak})$:
$\bm{Y_{a}}\sim\text{multinomial}(\sum_{k=1}^{K}x_{ak}\theta_{ak1},\dots,\sum_{k=1}^{K}x_{ak}\theta_{akK})$ (5)
Where $\theta_{akk^{\prime}}$ is the fraction of attention to a topic $k$ that has been redirected to a topic $k^{\prime}$. Moreover, we introduce intellectual capital $\bm{I_{a}}$ and social-intellectual capital $\bm{S_{a}}$ as covariates for transfers $\theta$ in the model, through a generalized linear model:
$\bm{\theta_{ak}}=\text{softmax}\left(\beta_{ak1}+\gamma_{k1}I_{a1}+\delta_{k1}S_{a1},\dots,\beta_{akK}+\gamma_{kK}I_{aK}+\delta_{kK}S_{aK}\right)$ (6)
[…]
Our model is strongly connected to Optimal Transport [49, 50]. Optimal Transport seeks the most efficient way to “transport” a source distribution (say, $\bm{x}=(x_{k})$) to a target distribution (e.g., $\bm{y}=(y_{k^{\prime}}$)) through “transfers” $(\theta_{kk^{\prime}})$ across their components, given a cost function $\sum_{k,k^{\prime}}x_{k}\theta_{kk^{\prime}}C_{kk^{\prime}}$ (where $C_{kk^{\prime}}$ is a cost matrix) [51]. The difference is that our Bayesian approach estimates transfers $\theta_{kk^{\prime}}$ by minimizing a likelihood rather than a cost function. The connection with Optimal Transport suggests an “economic” interpretation of the reallocation of research efforts, which will be leveraged in Section 3.2 to show that patterns of change in research interests are mainly structured by learning costs.
[…]
$\bm{x_{a}}$$\bm{y_{a}}$$\theta_{a}$$\beta_{a}$$\mu,\Sigma$$\bm{Z_{a}}$
Figure 2: 1(a),1(b),1(c): Physicists’ transfers of attention across research areas. $\bm{x_{a}}$ and $\bm{y_{a}}$ are the distributions of scientist $a$’s interests across research areas in two consecutive time periods. $\bm{\theta_{ak}}=(\theta_{akk^{\prime}})$ represents the fraction of the attention devoted by an author $a$ to a topic $k$ redirected to topics $k^{\prime}\in\{1,\dots,K\}$, as scientists repurpose their knowledge. By definition, $\sum_{k^{\prime}}\theta_{akk^{\prime}}=1$.
Figure 1(d): Hierarchical model. $\theta_{a}$ is drawn from a hierarchical process and also depends on a number of covariates $\bm{Z_{a}}$ (namely $\bm{I_{a}}$ and $\bm{S_{a}}$, intellectual and social “capital”). Observed variables are represented in gray, latent variables in white.
3 Results
3.1 Transfers across research areas
Figure 3 shows the aggregate transfers of attention at the level of the cohort unveiled by the model. The width of each flow is proportional to $\sum_{a}X_{ak}\theta_{akk^{\prime}}$ (summing over each author $a$). The most obvious feature is the overall conservatism of physicists toward their research area, i.e. specialization. Certain research areas have significantly grown, most notably “dark matter” and “AdS/CFT”, while others – “neutrinos and flavor physics” and “string theory and supergravity” – have significantly shrunk. The increasing interest in dark matter seems to be mainly fueled by a shift away from “neutrinos and flavor physics” and “electroweak sector”, while the growth in AdS/CFT and black hole research has been at the expense of “string theory and supergravity” research [^12] This converges with physicist Peter Woit’s controversial assessment that “string theorists” are no longer doing string theory per se, though they keep identifying themselves as string theorists. As Peter Woit puts it, citing the 2022 “Strings” conference: “one thing that stands out is that the string theory community has almost completely stopped doing string theory.”; and, “[presentations’ titles] make very clear what the string theory community has found to replace string theory: black holes” (Woit, 2022, https://www.math.columbia.edu/~woit/wordpress/?p=12981);. Smaller research areas such as “classical fields”, “dynamics” or “thermodynamics” seem less cohesive, i.e., they are subject to larger transfers. These bodies of knowledge are linguistically coherent (and therefore appear as topics of their own) but yet are not able to sustain careers over the long run by themselves.
Figure 3: Aggregate transfers of attention across research areas, between 2000-2009 (to the left) and 2015-2019 (to the right). Widths of flows are proportional to $\sum_{a}X_{ak}\theta_{akk^{\prime}}$. Insignificant transfers (that happen less than expected by chance alone assuming uniform mixing) are transparent.
It is insightful to look at specific careers. For instance, the two physicists who made the most abrupt switch from neutrinos to dark matter are shown in Figure 4. Physicist A’s personal website reads: “I am working on particle astrophysics and cosmology. In particular, I am interested in dark matter problem in the Universe, and how to probe it using annihilation products such as energetic gamma rays and neutrinos. […] I started my research career by studying supernova neutrinos from various aspects [our emphases].” [^13] https://staff.fnwi.uva.nl/s.ando/eng/Research.html. This example therefore constitutes an instance of “conversion” of prior knowledge to new purposes, one of the forms of change drawn from historical institutionalism represented in Figure 1. Similarly, Physicist B started his career in Neutrino Physics with a PhD on the “Semiclassical Analysis of Solar Neutrino Data”. His personal webpage reads: “My main research interests lie at the intersection of the fields of astrophysics, particle physics, and nuclear physics, concerning mostly neutrinos and the weak interactions. In the past several years, there has been strong growth at this intersection, in part due to new results from the ongoing experiments.” [^14] https://www.asc.ohio-state.edu/beacom.7/. Moreover, Physicist A claims that “Although [he is] a theoretical astrophysicist, [he] always consider prospects of current or near future generation of experiments and telescopes.”. This supports the assumption that experiments create opportunities (and pressures to adapt) that physicists can accommodate by building upon their prior knowledge.
(a) Physicist A.
(b) Physicist B.
Figure 4: Trajectory of the two physicists with the most significant shift from Neutrino physics to Dark matter. The initial distribution of the research agenda is shown in blue. The late distribution is shown in red. Rows represent research areas of origin and columns represent target research areas.
3.2 The structuring effect of learning costs on scientists’ behavior
As research priorities change, research efforts must be efficiently reallocated among scientists, given their prior knowledge. Here, we leverage the similarity between our model of scientists’ trajectories and Optimal Transport [49, 50] to formulate the observed behavior in economic terms and demonstrate that the shifts of attention are structured by the minimization of “learning costs”.
Let $\bm{x}=\sum_{a}\bm{x_{a}}$ be the cohort’s initial distribution of attention across research areas (summing over each author $a$), and $\bm{y}=\sum_{a}\bm{y_{a}}$ the late distribution. Let us further assume that $\bm{x}$ and $\bm{y}$ can be considered “fixed” by the institutions (scientific leadership, laboratories, funding agencies, etc.) that define scientific priorities throughout time. In order to achieve the distribution of research efforts $\bm{y}$ given the previous distribution $\bm{x}$, some scientists must redirect their attention away from certain research areas (for which $y_{k}<x_{k}$) and towards more pressing ones (for which $y_{k}>x_{k}$). What is the most efficient way to reallocate research efforts and achieve the transition from $\bm{x}$ to $\bm{y}$? Intuitively, research areas should be assigned to scientists in a way that requires as few of them as possible to acquire new knowledge – in other words, in a way that minimizes learning costs, given the way knowledge is distributed among individuals. This, as we show, can be framed as an Optimal Transport problem. Let $T_{kk^{\prime}}$ be the coupling matrix that encodes how much attention has been shifted from one research area $k$ to a research area $k^{\prime}$. There are two constraints on $T_{kk^{\prime}}$: $\sum_{k^{\prime}}T_{kk^{\prime}}=x_{k}$, and $\sum_{k}T_{kk^{\prime}}=y_{k^{\prime}}$. These constraints encode the need to adapt (since $\bm{y}\neq\bm{x}$). But $T$ must also minimize learning costs ($C$), which are linear in $T$: $C=\sum_{k,k^{\prime}}T_{kk^{\prime}}C_{kk^{\prime}}$, where $C_{kk^{\prime}}$ is a cost matrix. The problem of finding the couplings $T_{kk^{\prime}}$ that minimize the “transportation” costs (given a cost matrix $C_{kk^{\prime}}$ and the constraints on $T_{kk^{\prime}}$) is an instance of Optimal Transport problem [51]. Examples of Optimal Transport problems include how to efficiently transport ore from a set of mines to a set of factories given the distances between them, or the optimal assignment of workers to firms [52]. In our case, the couplings are known (they have been recovered by our model from Section 2.4), and we would like to infer the underlying cost matrix that these couplings minimize. The transfers from a research area $k$ to $k^{\prime}$ for each individual are simply $x_{ak}\theta_{akk^{\prime}}$. Summing over individuals yields the coupling matrix $T_{kk^{\prime}}=\sum_{a}x_{ak}\theta_{akk^{\prime}}$, which measures how much attention was shifted away from $k$ and toward $k^{\prime}$ at the cohort level.
[…]
Under changing circumstances, research efforts must be reallocated efficiently. Therefore, scientific norms and institutions address an Optimal Transport problem by providing incentives for scientists to conform to new research imperatives, in a way that minimizes learning costs. While this observation characterizes the cohort behavior, drivers of differences among individuals are considered next.
3.3 Individual behavior and the effect of capital
3.3.1 Magnitude of change and capital
We propose a change score $c_{a}$ that measures how much the research agenda of a scientist has changed between the two time periods under consideration, defined for each physicist as the total variation distance between their initial and late research agendas:
$c_{a}\equiv\mathrm{d}_{\text{TV}}(\bm{x_{a}},\bm{y_{a}})=\frac{1}{2}\sum_{k}|y_{ak}-x_{ak}|$ (12)
$c_{a}$ is comprised between 0 (if research attention has remained identically distributed) and 1 (if the research agenda has been entirely redistributed). The distribution of $c_{a}$ (Figure 5) shows that large values are rare, with 50% of authors lying between 0.21 and 0.40. However, the absolute value of $c_{a}$ has limited interpretability, given that it depends on the choice of “cognitive scope” (the broadness of research areas) and the duration of the time periods considered [13]. Moreover, $c_{a}$ is never expected to be exactly zero due to random fluctuations.
Figure 5: Distribution of change scores in the cohort. Higher values correspond to more drastic changes in a scientists’ research agenda.
Nevertheless, the change score allows for comparisons between physicists. We evaluate the effect of several factors on change: i) the diversity of intellectual capital $D(\bm{I_{a}})$; ii) the excess diversity of social capital $D^{\ast}(\bm{S_{a}})$; and iii) the magnitude of social capital (“power”). We also consider the effect of affiliation stability, represented by a binary variable $s_{a}$ ($s_{a}=1$ if scientist $a$ has at least one affiliation that spans the whole time range considered, and $s_{a}=0$ otherwise). We perform a linear regression of $c_{a}$ as a function of these four variables, adjusting for $Z_{a}=\operatorname*{arg\,max}_{k}x_{ak}$, i.e. physicists’ primary research area over the years 2000 to 2009 (see the model specification in Appendix A.7.1).
The results are shown in Figure 6. The diversity of intellectual capital has a significant positive effect: physicists with resources in many areas tend to revise their research agenda more. There is also evidence of a positive effect of diversity of social capital on the magnitude of changes in scientists’ research focus. However, the magnitude of social capital, “power”, has a negative direct effect on change. In other words, “power” promotes stability, while “diversity” promotes change. It is noteworthy that these dimensions of social capital have opposite effects. Physicists with permanent affiliations are more conservative toward their research agenda, possibly because they experience less incentive to “adapt”. Conversely, new positions might be associated with breaks in research topics.
Unsurprisingly, both research areas that have shrunk considerably have a significant positive effect on migration score. “Collider physics”, on the other hand, has a negative effect on the magnitude of change. This is the case ceteris paribus; in other words, physicists whose primary category is “Collider physics” are more conservative than other physicists with similar capital diversity, power, and position stability. In fact, these physicists also have less diverse capital, and are by far the most conservative, with an average change score 24% lower than the rest of the cohort. Indeed, the long time-scales of collider experiments provide stable opportunities to physicists in that area [55, p. 138], thus promoting conservatism. The variance explained remains low ($R^{2}=0.15$), and all these factors only partially explain differences between individuals.
Figure 6: Effect of measures of intellectual and social capital, and of position stability on the amplitude of changes in high-energy physicists’ interests (as measured by the change score $c_{a}$). Diversity is evaluated in terms of the exponential of the Shannon entropy of the distribution of intellectual and social resources physicists possess across research areas. The effect of the primary research area during the initial period (2000-2009) was also included as a control. Three research areas have a significant (95% CL) effect. Error bars indicate 95% credible intervals; $\mu(\beta)$ denotes the mean posterior effect size. $R^{2}=0.15$. Continuous variables (diversity, power and the change score) are standardized.
The change score $c_{a}$ neglects the cognitive gap between research areas. Figure 15 in Appendix A.7.3 shows that, using an alternative measure of change (“disruption”) that takes into account cognitive distance, the diversity of social capital has a much larger effect ($\mu(\beta)=0.28$ instead of $0.091$). Moreover, using the Stirling index (rather than entropy) as a measure of diversity does not qualitatively change the previous results (see Table 3, Appendix A.7.3). Using the Stirling index as a measure of diversity, the diversity of social capital has an even stronger effect on disruption ($\mu(\beta)=0.85$).
3.3.2 Diversification of research interests versus consolidation of expertise
Research portfolios can be altered by two opposite strategies. One strategy is “diversification”, through the addition of new research areas. Another strategy is “concentration”, by deserting certain research areas and concentrating increasingly on a core specialization (Figure 1 illustrates how this can happen, whether via “conversion” or “displacement”). Figure 7 shows the effect of the same factors as above on the probability that physicists have i) entered at least one new research area in between the two periods or ii) exited one research area (see the specification of the model in Appendix A.7.2) [^16] A research area $k$ is considered “entered” by $a$ when $x_{ak}<\frac{1}{N}\sum_{a^{\prime}}x_{a^{\prime}k}$ and $y_{ak}>\frac{1}{N}\sum_{a^{\prime}}y_{a^{\prime}k}$; conversely, a research area is considered exited when $x_{ak}>\frac{1}{N}\sum_{a^{\prime}}x_{a^{\prime}k}$ and $y_{ak}<\frac{1}{N}\sum_{a^{\prime}}y_{a^{\prime}k}$..
The diversity of intellectual capital has a strong positive effect on the probability of exiting a research area; intuitively, scientists with diverse interests can afford to disengage from certain research areas. Excess diversity of social capital increases the probability of entering new research areas, but has no discernible effect on the probability of exiting research areas. In contrast, “power” significantly decreases the probability of leaving a research area. This suggests that having many collaborators allows scientists to remain committed to many research areas with minimal personal investment, thus stabilizing their research agendas (as previously observed). Moreover, “power” also decreases the probability to of entering new research areas; this suggests that powerful scientists have less incentives to invest resources in new topics. String theory & supergravity has a significant positive effect of both entering and exiting research areas (suggesting it has been “replaced” by another topic by many physicists). “Collider physics”, however, significantly decreases the probability of entering new areas. Indeed, the LHC has secured this research area for the decades to come: for instance, the HL-LHC project (High-Luminosity LHC), which consists in upgrading the LHC in order to higher collision rates, is scheduled to start in 2029 at the time of writing [^17] See https://home.cern/science/accelerators/high-luminosity-lhc. Note that until a successor to the LHC is under active development, its prolongation can help preserve the know-how necessary for the design and operation of such complex experiments.. “Dark matter” has a negative effect on the probability of entering a new research area, possibly because scientists with prior commitment to this research area had less incentive to diversify their research portfolio and more incentive to focus increasingly on this topic given its success over the following years. Again, these conclusions hold as an alternative measure of diversity is considered (A.7.3, Table 4) and if “brokerage” is used as a measure of power.
Figure 7: Effect of capital and position stability on the probability of entering new research areas or exiting previously explored ones. Bars indicate 95% credible intervals. $\mu(\beta)$ denotes the mean posterior effect size. Continuous variables (diversity and power) are standardized.
3.3.3 Why diversity promotes change
Access to diverse cognitive resources promotes change. To see why, it is insightful to look into how the concentration of scientists’ intellectual capital in each research area $k$ ($\bm{I_{a}}=(I_{ak})$) affects their trajectories. In the model introduced in Section 2.4, the diagonal coefficients of the $\gamma$ matrix measure the effect of having intellectual resources in a certain area on the commitment to this area. The matrix is shown in Figure 7(a). Most coefficients on the diagonal are significantly positive, thus suggesting that physicists with a strong specialization in a research area have tend to further exploit and consolidate their specialization in this area.
The effect of social capital on transfers from one research area to another, given by the matrix $\delta_{kk^{\prime}}$ (see equation 6), is shown in Figure 7(b). Statistically significant effects are always positive: having collaborators with expertise in a target area generally increases the likelihood of redirecting attention toward this area.
(a) Effect of intellectual capital ($\gamma_{kk^{\prime}}$).
(b) Effect of social capital ($\delta_{kk^{\prime}}$).
[…]
The effect of social capital varies across research areas. Indeed, [7, ] found that “social-relatedness” (i.e. whether scientists have immediate access to a subfield via a collaborator) is more crucial as scientists move further away from their specialization, and our model reproduces this finding. The effect of social capital ($\delta_{kk^{\prime}}$) depends on the matrix $\nu_{kk^{\prime}}\in[0,1]$, defined as the fraction of physicists with more expertise than average in a research area $k^{\prime}$ among those who have more expertise than average in $k$ (see Eq. (9), Section 2.4). We find a strong association, with $\delta_{kk^{\prime}}$ decreasing by 1.5 unit of standard deviation on average as $\nu_{kk^{\prime}}$ goes from 0 (nobody holds expertise in both research areas $k$ and $k^{\prime}$) to 1 (everyone with expertise in $k$ has expertise in $k^{\prime}$). In other words, the strength of the effect of social capital on shifts between two areas is higher when it is uncommon to hold expertise in both these areas.
4 Discussion
This paper addressed the conflict between specialization and conservatism in science, which are driven by the need to fully exploit accumulated knowledge, and the urge to adapt to changes in the epistemic landscape. To this end, an unsupervised Bayesian approach was developed, based on the intuition that scientists’ shifts in research interests involve cumulative transfers of knowledge. The model simultaneously measures transfers of attention across research areas and the effect of various variables on the evolution of scientists’ research portfolios, in particular intellectual and social capital.
[…]
Then, leveraging the connection between our model and Optimal Transport (OT), we showed that the reallocation of research efforts among scientists is structured by learning costs. Indeed, under changing circumstances, scientific institutions must address an OT problem by efficiently reallocating research efforts in a way that balances learning costs and the imperative to adapt to new circumstances. This has the effect of enhancing the utility of the scientific capital disseminated amongst scientists as the perceived payoff of certain research areas increases. The connection with OT is methodologically useful: Inverse Optimal Transport allows one to derive cost functions from observed behavior, and thus offers a potential way to better connect empirical data with formal rational-choice models of scientists’ behavior. Moreover, it has the potential of informing policy by identifying potential bottlenecks if research efforts were to be reallocated in certain ways (as one can use OT to estimate the “cost” of various counterfactual scenarios).
The longitudinal comparative analysis of physicists’ trajectories revealed that the diversity of intellectual and social capital is positively associated with change: diversity promotes adaptability under new circumstances, and therefore diversifying research portfolios is a reasonable strategy when the future is uncertain. However, enforcing diversity can lead to suboptimal allocations of research efforts under stable circumstances [25]. Differences among research areas are found: physicists expert in particle colliders have remained particularly conservative, possibly because they have secured long-term research opportunities (thanks to very large investments in particle accelerators like the LHC). There is also evidence that physicists specialized in dark matter have been consolidating their specialization, presumably because the increasing popularity of the topic encouraged them to double down their investment in this research area. Higher concentrations of intellectual capital in certain research areas generate stronger commitment to these research areas; therefore, specialized scientists are more at risk of being trapped in a sunk cost fallacy as their expertise becomes unsuitable for new circumstances. However, specialized scientists can offset the risks associated with specialization by diversifying their social network. This raises the possibility of free-riding, as scientists are encouraged to focus on what seems most promising at the time and let their peers take the risk of exploring alternatives until their value is established [56]. Additionally, social capital plays an increasingly important role as scientists expand their research agenda further beyond their specialization, as observed by [7], suggesting that collaborations are crucial in overcoming cognitive barriers between research areas. Unlike diversity, “power” (that is, the magnitude of social capital) and institutional stability are associated with more stability and conservatism in research interests. We suggest that coalitions and institutional resources can safeguard individuals from pressures to adapt. The negative association between change and institutional stability, albeit weak ($\mu(\beta)=-0.14$), suggests that scientific institutions may exploit temporary contracts to compell the scientific workforce to adapt.
We have characterized renewal strategies of research portfolios in terms of a typology of incremental change developed in the context of historical institutionalism, which has been shown to account for how organizations like DESY and SLAC have transitioned from HEP to multi-purpose photon science. Adaptation strategies include the “conversion” of prior knowledge to new purposes, the “layering” of additional research interests via the acquisition of new knowledge, or the “displacement” of former research commitments. Indeed, institutions can face a challenge similar to that experienced by specialized scientists who are confronted with transformations in their scientific landscape, in that they too must adapt and redirect accumulated capital in directions that may not have been foreseen at the time of their “design”. We conclude that a better understanding of adaptation benefits from the pooling of diverse insights: while studies of institutional change have documented strategies of adaptation when change is difficult, the literature on cultural evolvability stresses the critical roles of “diversity” and social learning. In return, this work provides an empirical contribution to the literature that treats science as a cultural evolutionary system [12].
Throughout the paper, various methodological limitations were raised. Most of them pertain to data selection. For example, the requirement on the quantity of authors’ publications makes the cohort “unrepresentative” and implies a survivorship bias. A second issue, already noted by [13, ], lies in the arbitrariness in the choice of cognitive scope. Previous works (e.g. [4, 5, 14, 7]) have typically relied on the hierarchical PACS classification of physics literature, such that cognitive scales were imposed. For this work, the literature was divided into 16 topics that captured the features of change in the epistemic landscape discussed above, i.e. the rise of new probes of the cosmos, but other scales could have been considered. However, the topic model recovers a flat structure that does not explain the correlations between different research areas. A hierarchical structure of knowledge might be more suitable for capturing what allows physicists to switch from one research area to another.
Acknowledgements
I thank Thomas Heinze, Radin Dardashti, Elisa Omodei, and Ulysse Lojkine for useful discussions; Georges Ricci for his invitation to look into Optimal Transport; the LATTICE lab in Montrouge for the opportunity to exchange about this work; and Elizabeth Zanghi for corrections.
Funding
This work has received support from the G-Research Grant for Early-Career Researchers.
Use of Artificial Intelligence
Artificial Intelligence was not used in the process of writing this paper.
Data, code and materials
The data and code used for this publication will be made available in a public repository.
Competing Interests
The author declares no competing interest.
Appendix A Supplementary material
A.1 Transformations in high-energy physics
Figure 9: Share of citations of each of the most cited experiments in high-energy physics literature, between 2000 and 2019. Hatched rectangles correspond to experiments observing particles produced in colliders or nuclear reactors; other rectangles correspond to observations of phenomena or particles of astrophysical origin. We use citation and experiment data from the Inspire HEP database [36]. The classification of these experiments is our own.
A.2 Cohort characteristics
Figure 10: Distribution of academic age (in 2015) and affiliation stability within the considered cohort. Academic age (left plot) is evaluated as the time passed since the first publication. There is a cut-off in the distribution academic age due to the requirement on publication counts between 2000 and 2009. The median age is 23 years in 2015. 49% of the cohort has had at least one permanent affiliation spanning the period under consideration (right plot). As expected, permanent affiliations are associated with higher academic ages (bars indicate min/max/quartiles).
A.3 Topics
A.3.1 Keyword classification
In the Embedding Topic Model used in the present paper – as in topic models in general–, documents are mixtures of several topics; moreover, keywords may belong to several topics. This is a desirable feature: certain concepts serve different purposes depending on the context, and some concepts do not clearly belong to any research area. More importantly, this is crucial to the present approach, given that we must be sensitive to situations in which scientists have repurposed certain concepts to new research goals; i.e., instances where the same resources are applied in a new context (a new research area). For that, each keyword $i$ from a document $d$ is assigned a research area $z_{di}:=\operatorname*{arg\,max}_{k=1,\dots,K}P(z_{di}=k|w_{di},\bm{\theta}_{d})$, which is the research area most probably associated with the keyword $i$ given the topic distribution of the document $\bm{\theta}_{d}$ – i.e. the context. In fact:
$z_{di}:=\operatorname*{arg\,max}_{k=1,\dots,K}P(z_{di}=k|w_{di},\bm{\theta}_{d})=\operatorname*{arg\,max}_{k=1,\dots,K}\dfrac{P(w_{di}|z_{di}=k)P(z_{di}=k|\bm{\theta}_{d})}{P(w_{di}|\bm{\theta}_{d})}$ (13)
In the process, we discard ambiguous keywords for which $H(z_{di})\geq\log{2}$, where $H$ denotes the entropy of the distribution $P(z_{di}=k|w_{di})$. Either such keywords do not carry any scientific content, or the context is insufficient to disambiguate among the possible research areas to which they might belong.
The average effective amount of topics in documents according to the topic model – measured as $\exp H(\bm{\theta}_{d})$ where $H$ is the Shannon entropy – is 7.5, which is unrealistically high. By performing this filtering and classification procedure, the average effective amount of topics goes down to 3.2 per document, which is much more informative and realistic. This extra step is especially important for short texts such as abstracts, which poorly constrain the latent topic distribution $\bm{\theta}_{d}$.
A.3.2 List of topics
Table 1: Research areas, their top-words, and their correlation with a standard classification (PACS).
Research area
Top words
Most correlated PACS categories
AdS/CFT
boundary, holographic, flow, bulk, critical, conformal, critical_point, boundary_theory, cfts, point, bootstrap, central, conformal_anomaly, strip, fixed, free, entanglement_entropy, conformal_field_theory, criticality, condition
Gauge/string duality (0.27) Conformal field theory, algebraic […] (0.23) Theory of quantized fields (0.15) Theories and models of […] (0.14) Critical point phenomena (0.14)
Black holes
hole, black_hole, gravity, black, horizon, geometry, gravitational, spacetimes, spacetime, curvature, thermodynamics, einstein, schwarzschild, metric, ad, hair, relativity, space_time, observer, graviton
Quantum aspects of black holes, […] (0.51) Classical black holes (0.41) Physics of black holes (0.32) Exact solutions (0.24) Higher-dimensional black holes, […] (0.23)
Classical fields
scalar, general, first, scalar_field, massless, real, explicit, dynamical, second, exact, special, linear, full, symmetric, static, electromagnetic, classical, nonlinear, approximate, non_trivial
Modified theories of gravity (0.19) Lorentz and Poincaré invariance (0.16) Nonlinear or nonlocal theories and […] (0.11) Exact solutions (0.10) Higher-dimensional gravity and […] (0.10)
Collider physics
distribution, collision, production, cross_sections, section, parton, hadron, cross, cross_section, process, hadronic_collision, scattering, correction, fragmentation, partons, kinematics, transverse, impact, event, partonic
Perturbative calculations (0.29) Polarization in interactions and […] (0.28) Inclusive production with […] (0.27) Total and inclusive cross sections […] (0.26) Relativistic heavy-ion collisions (0.25)
Cosmology
constant, cosmological, inflation, cosmic, perturbation, vacuum, universe, inflationary, cosmology, fluctuation, inhomogeneity, tension, lambda, planck, inflaton, cosmological_perturbation, era, epoch, density, background
Particle-theory and field-theory […] (0.59) Cosmology (0.32) Observational cosmology (including […] (0.28) Dark energy (0.25) Background radiations (0.21)
Dark matter
matter, dark, dark_matter, detection, dm, signal, abundance, observation, relic, direct, constraint, candidate, wimp, asymmetric, prospect, dark_matter_particle, center, detectable, cold, contribute
Dark matter (0.74) γ-ray (0.22) Cosmic rays (0.19) γ-ray sources; γ-ray bursts (0.17) Elementary particle processes (0.17)
Dynamics
solution, equation, phase, space, time, system, transition, region, condition, constraint, dynamic, class, background, configuration, wave, range, motion, set, star, instability
Exact solutions (0.14) Nonlinear or nonlocal theories and […] (0.11) Extended classical solutions; […] (0.10) Relativistic wave equations (0.10) Modified theories of gravity (0.09)
Electroweak sector
standard, higgs, boson, particle, standard_model, physic, lhc, sm, top, tev, collider, mssm, electroweak, minimal, phenomenology, extension, extra, supersymmetric_model, superpartners, new_particle
Extensions of electroweak Higgs sector (0.34) Supersymmetric models (0.33) Non-standard-model Higgs bosons (0.30) Supersymmetric partners of known […] (0.28) Standard-model Higgs bosons (0.27)
Gauge theory
dimension, coupling, scale, structure, operator, fermion, value, matrix, number, su, charge, sector, spin, group, topological, anomalous, breaking, anomaly, global, flavor
Unified theories and models of […] (0.22) Unification of couplings; mass relations (0.17) Quark and lepton masses and mixing (0.13) Unified field theories and models (0.12) Field theories in dimensions other […] (0.12)
Hadrons
decay, data, channel, bound, resonance, gamma, meson, width, experimental_data, collaboration, kaon, prediction, experimental, measurement, admixture, narrow, process, hadronic_decay, s0, ratio
Decays of bottom mesons (0.30) Decays of J/ψ, Υ, and other quarkonia (0.24) Meson-meson interactions (0.21) Decays of bottom mesons (0.20) Bottom mesons (|B|>0) (0.19)
Neutrinos and flavour physics
neutrino, violation, oscillation, flavor, cp, angle, mixing, experiment, lepton, flavour, hierarchy, majorana, cp_violation, beta, leptogenesis, asymmetry, neutrino_mass, neutrino_oscillation, smallness, generation
Neutrino mass and mixing (0.74) Non-standard-model neutrinos, […] (0.41) Ordinary neutrinos (0.30) Neutrino interactions (0.28) Quark and lepton masses and mixing (0.23)
Perturbative methods
amplitude, qcd, loop, diagram, sum, contribution, perturbative, expansion, vertex, rule, light_cone, perturbative_qcd, propagator, approach, correlator, one_loop, evaluation, nonperturbative, kernel, diagrammatic
General properties of perturbation […] (0.25) Other nonperturbative calculations (0.24) Sum rules (0.22) Perturbative calculations (0.21) General properties of QCD […] (0.16)
QCD
quark, chiral, magnetic, baryon, relativistic, moment, qcd, light_quark, strong, heavy, heavy_quark, lattice, magnetic_field, electric, deconfinement, chromodynamics, current, diquarks, plasma, color
Lattice QCD calculations (0.27) Chiral symmetries (0.26) Chiral Lagrangians (0.25) Quark-gluon plasma (0.23) General properties of QCD […] (0.20)
Quantum Field Theory
quantum, group, quantum_field, representation, quantisation, mechanic, quantum_field_theory, transformation, hamiltonians, algebra, finite_dimensional, quantization, commutator, algebraic, arbitrary, operator, qft, invariant, analog, associated
Algebraic methods (0.26) Noncommutative field theory (0.25) Quantum mechanics (0.22) Noncommutative geometry (0.19) Quantum groups (0.18)
String theory and supergravity
string, supersymmetric, superstring, six_dimensional, modulus, super, instantons, supergravity, dyons, n2, mathcaln, superpotentials, heterotic, sigma_models, n1, n4, gauged, space, deformation, compactifications
Supersymmetry (0.31) Strings and branes (0.29) Supergravity (0.29) Compactification and four- […] (0.25) D branes (0.20)
Thermodynamics
potential, effective, interaction, limit, temperature, action, finite, local, freedom, approximation, level, weak, chemical, force, effective_field_theory, lagrangian, finite_temperature, effective_field, degree, effective_theory
Finite-temperature field theory (0.26) Chiral symmetries (0.09) Nuclear forces (0.08) Quark-gluon plasma (0.08) General properties of QCD […] (0.08)
Uninterpretable
approach, method, analysis, recent, calculation, numerical, formalism, study, prediction, sigma, previous, work, theoretical, systematic, comparison, uncertainty, agreement, good, investigation, paper
Lattice QCD calculations (0.07) Baryon resonances (S=C=B=0) (0.05) Other nonperturbative calculations (0.05) Few-body systems (0.05) Lagrangian and Hamiltonian approach (0.05)
Uninterpretable
form, correction, momentum, tensor, mode, relation, higher, factor, vector, invariant, formula, angular, part, theorem, spectrum, power, dimensional, invariance, expression, derivative
Electromagnetic form factors (0.17) Protons and neutrons (0.10) Lorentz and Poincaré invariance (0.08) Gauge field theories (0.06) Dispersion relations (0.06)
Uninterpretable
spectrum, low, problem, low_energy, important, high, property, high_energy, small, physical, soft, fundamental, behavior, analytic, behaviour, spectral, dispersion, essential, phenomenon, regime
General properties of QCD […] (0.07) Regge formalism (0.07) Wave propagation, transmission and […] (0.05) Elastic scattering (0.05) Lattice gauge theory (0.05)
Uninterpretable
different, possible, particular, present, various, mechanism, type, example, massive, several, scenario, simple, single, similar, consistent, addition, hand, different_type, interesting, way
Particle-theory and field-theory […] (0.07) Modified theories of gravity (0.07) Field theories in dimensions other […] (0.05) Cosmology (0.05) Dark energy (0.05)
A.3.3 Topic validation using the citation network
In order to validate the consistency and scientific dimension of the topics that were recovered, we verify that papers from a given topic tend to cite more papers from the same topic. Let $N_{k,k^{\prime}}$ be the amount of citations of articles that belong to topic $k^{\prime}$ originating from articles that belong to $k$, and $N=\sum_{k,k^{\prime}}N_{k,k^{\prime}}$ the total number of citations. From this matrix, a normalized pointwise mutual correlation $\mathrm{npmi}(k,k^{\prime})$ is calculated:
$\mathrm{npmi}(k,k^{\prime})=\log{\dfrac{N_{k,k^{\prime}}/N}{(\sum_{i}N_{k,i}/N)(\sum_{i}N_{i,k^{\prime}}/N)}}$ (14)
$\mathrm{npmi}(k,k^{\prime})$ is shown in Figure 11. It measures how frequent citations from $k$ to $k^{\prime}$ are, relative to what would be expected if citations were uniformly distributed. The diagonal values are positive, indicating that the topics we retrieved tend to refer to themselves significantly more than expected by chance alone, providing further evidence of their scientific content and coherence.
Figure 11: Normalized pointwise mutual information $\mathrm{npmi}(k,k^{\prime})$ of the citation matrix $N_{k,k^{\prime}}$. Positive values (in blue) indicate that the y-axis research area cites the x-axis research area more often than expected by chance; negative values indicate that citations occur less than expected by chance. When shown, individual values indicate how many times citations occur compared to chance alone. For instance, papers about Hadrons cite papers about Perturbative methods 1.4 times what would be expected if citations were uniform across research areas.
A.3.4 Topic validation using the PACS classification
Figure 12: Correlation between PACS categories present in $\geq 100$ publications in the corpus (rows), and the research areas recovered by the topic model (columns).
A.4 Measures of capital
Figure 13 shows the Pearson correlation between these diversity and power measures, as evaluated in the cohort of high-energy physicists. For comparison purposes, an alternative measure of diversity based on Stirling’s index [57], with prior applications to studies of interdisciplinarity [58, 59] is evaluated [^18] The Stirling-based diversity measured is evaluated as: $D_{\text{Stirling}}=1-\sum_{k,k^{\prime}}d_{kk^{\prime}}I_{ak}I_{ak^{\prime}}$ (15) Where $d_{kk^{\prime}}$ is the fraction of scientists who have more expertise than average in both $k$ and $k^{\prime}$ among those that have expertise in one or the other (i.e., a similarity matrix); see [58, 59]. A measure of brokerage is also considered [^19] We evaluated brokerage as the amount of pairs of scientists that have collaborated with a given physicist while having no common collaborator except for this physicist. This effectively measures the extent to which this physicist connects otherwise disconnected scientists..
As shown in Figure 13, the entropic measure of diversity considered in this paper correlates strongly with the Stirling measure. The magnitude of social capital (which is similar to degree centrality) correlates weakly with excess social diversity, thus emphasizing that power and diversity are partially orthogonal aspects of social capital. Power is strongly correlated with brokerage; indeed, strongly connected scientists, with higher degree centrality, are also those scientists who initiate collaborations between otherwise disconnected scientists, as measured by brokerage. Most importantly, brokerage correlates slightly better with measures of intellectual diversity than with excess social diversity. This suggests that the benefits of “bridging” positions in social networks can sometimes be confused with the benefits of the actors’ own diversity of intellectual resources, which enables them to occupy such positions in the first place. This emphasizes the value of distinguishing the effect of individual intellectual capital from that of social capital.
Figure 13: Correlation between different measures of capital. Measures considered in this analysis are shown in bold. Alternative measures are shown for comparison purposes. By construction, excess social diversity is orthogonal to intellectual diversity.
A.5 Model performance
The predictive power of the model can be assessed by evaluating the total variation distance between the true distribution $\bm{y_{a}}$ and the predicted distribution $\bm{y_{a}^{\text{pred}}}$. This performance metric is calculated via 10-fold cross-validation. It is compared to a baseline model that predicts no change in the research agenda ($\bm{y_{a}^{\text{baseline}}}=\bm{x_{a}}$). The results are shown in Table 2. For the cohort of high-energy physicists, the model performs only marginally better than the baseline, given that individuals have remained quite conservative on average, most of the fluctuations being difficult to predict. Table 2 also considers a cohort of scientists from the ACL anthology corpus of computational linguistics research [15], by running the same pipeline (the measurement of research portfolios during two consecutive time periods using our topic model approach and the training of the trajectory model). Although the data are of lesser quality (see Appendix A.9, Table 7), research portfolios have undergone much more significant transformations in this dataset (see Appendix A.9, Figure 16). Consequently, our model performs much better than the baseline for this cohort.
Table 2: Performance of the actual model versus that of the baseline model, for i) the cohort of high-energy physicists and ii) a cohort from the ACL Anthology corpus of Computation Linguistics papers.
Cohort
High-energy physics
($N=2195$)
Computational Linguistics
($N=578$)
Model
$\mu(d_{{\mathrm{{TV}}}}(\bm{{y_{{a}}}},\bm{{y_{{a}}}}^{{\text{{pred}}}}))$
0.306 0.337
Baseline
$\mu(d_{{\mathrm{{TV}}}}(\bm{{y_{{a}}}},\bm{{x_{{a}}}}))$
0.316
0.466
A.6 Learning costs and optimal transport
The MCMC algorithm from [53] is run on 500,000 iterations. The mean posterior values of $C_{k,k^{\prime}}$ are shown in Figure 13(a), as a function of the linguistic gap from $k$ to $k^{\prime}$, $\Delta_{kk^{\prime}}=H(\bm{\varphi_{k^{\prime}}}+\bm{\varphi_{k}})-H(\bm{\varphi_{k}})$ (where $H$ is the Shannon entropy and $\bm{\varphi_{k}},\bm{\varphi_{k^{\prime}}}$ are the vocabulary distributions of $k$ and $k^{\prime}$, such that $\varphi_{kw}$ is the frequency of the keyword $w$ in the topic $k$). For replication purposes, the same procedure is performed on a different dataset – the ACL anthology corpus of computational linguistics research [15]. The correlation is smaller for the ACL cohort ($R=0.82$ instead of $0.96$). The difference could be explained by a number of reasons: a) the smaller sample size of the ACL cohort ($N=578$ instead of $N=2195$); b) differences in the quality of the topic models c) the cognitive scale (on which the overlap in vocabulary between research areas depends) is not commensurable across the two corpora, and finally d) the relationship does not hold as well for fields undergoing major disruptions.
(a) High-energy physics
(b) Computational Linguistics
Figure 14: Cost of shifting a unit of cohort’s research attention from $k$ to $k^{\prime}$ as a function of the fraction of physicists with expertise in $k^{\prime}$ among those with expertise in $k$ ($\nu_{kk^{\prime}}$).
A.7 Effect of capital on strategies of change
A.7.1 Model for the magnitude of change
The model for $c_{a}$ is:
$z(c_{a})$ $\sim\mathcal{N}(\mu_{a},\sigma)$
$\mu_{a}$ $=\beta^{\text{int-div}}z(D(\bm{I_{a}}))+\beta^{\text{soc-div}}z(D^{\ast}(\bm{S_{a}}))+\beta^{\text{power}}z(P(\bm{S_{a}}))+\beta^{\text{stability}}p_{a}+\mu^{\text{area}}_{k_{a}}+\mu$
$k_{a}$ $=\operatorname*{arg\,max}_{k}x_{ak}$
$\beta,\mu$ $\sim\mathcal{N}(0,1)$
$|\mu^{\text{area}}_{k}|$ $\sim\mathrm{Exponential}(\tau)$
$\tau,\sigma$ $\sim\mathrm{Exponential}(1)$
Where $z(\cdot)$ denotes standardized variables.
A.7.2 Model for the probability of having entered/exited a research area
The model for the probability $p_{a}$ of having entered a new field:
$\mathrm{logit}(p_{a})$ $=\beta^{\text{int-div}}z(D(\bm{I_{a}}))+\beta^{\text{soc-div}}z(D^{\ast}(\bm{S_{a}}))+\beta^{\text{power}}z(P(\bm{S_{a}}))+\beta^{\text{stability}}p_{a}+\mu^{\text{area}}_{k_{a}}+\mu$
$k_{a}$ $=\operatorname*{arg\,max}_{k}x_{ak}$
$\beta,\mu$ $\sim\mathcal{N}(0,1)$
$|\mu^{\text{area}}_{k}|$ $\sim\mathrm{Exponential}(\tau)$
$\tau,\sigma$ $\sim\mathrm{Exponential}(1)$
The same model structure is used for the probability of having exited a research area.
A.7.3 Effect of capital and robustness checks
Figure 15: Effect of standardized measures of intellectual and social capital, and position stability on the disruption of the physicists’ research agenda (as measured by the disruption score $d_{a}$). The disruption score is defined as the minimum cost of “transporting” $\bm{x_{a}}$ to $\bm{y_{a}}$ given the cost matrix $C_{kk^{\prime}}=1-\nu_{kk^{\prime}}$ (see Section 3.2). Disruption is higher for scientists who migrate to cognitively distant research areas.
Table 3: Effect of each variable on (a) the change score and (b) the disruption score for each model. The reference model uses entropy as the diversity measure $D$ and the magnitude of intellectual capital as a measure of power $P$. Values indicate the mean posterior effect size and the 95% credible interval. Significant effects are shown in bold.
Dep. variable Change score ($c_{a}$) Disruption score ($d_{a}$)
Model Reference $D=\text{Stirling}$ $P=\text{Brokerage}$ Reference $D=\text{Stirling}$ $P=\text{Brokerage}$
Predictor
Intellectual capital (diversity) $\bm{+0.28}\begin{subarray}{c}+0.043\\ -0.043\end{subarray}$ $\bm{+0.28}\begin{subarray}{c}+0.041\\ -0.041\end{subarray}$ $\bm{+0.22}\begin{subarray}{c}+0.043\\ -0.043\end{subarray}$ $\bm{+0.04}\begin{subarray}{c}+0.04\\ -0.04\end{subarray}$ $\bm{+1.2}\begin{subarray}{c}+0.47\\ -0.48\end{subarray}$ $+0.04\begin{subarray}{c}+0.04\\ -0.04\end{subarray}$
Social capital (diversity) $\bm{+0.09}\begin{subarray}{c}+0.04\\ -0.04\end{subarray}$ $\bm{+0.07}\begin{subarray}{c}+0.042\\ -0.041\end{subarray}$ $\bm{+0.06}\begin{subarray}{c}+0.04\\ -0.04\end{subarray}$ $\bm{+0.28}\begin{subarray}{c}+0.045\\ -0.044\end{subarray}$ $\bm{+0.85}\begin{subarray}{c}+0.47\\ -0.48\end{subarray}$ $\bm{+0.22}\begin{subarray}{c}+0.043\\ -0.044\end{subarray}$
Social capital (power) $\bm{-0.22}\begin{subarray}{c}+0.041\\ -0.041\end{subarray}$ $\bm{-0.2}\begin{subarray}{c}+0.04\\ -0.04\end{subarray}$ $\bm{-0.14}\begin{subarray}{c}+0.044\\ -0.044\end{subarray}$ $\bm{-0.23}\begin{subarray}{c}+0.042\\ -0.042\end{subarray}$ $\bm{-0.21}\begin{subarray}{c}+0.04\\ -0.041\end{subarray}$ $\bm{-0.14}\begin{subarray}{c}+0.044\\ -0.045\end{subarray}$
Stable affiliation $\bm{-0.14}\begin{subarray}{c}+0.079\\ -0.08\end{subarray}$ $\bm{-0.14}\begin{subarray}{c}+0.079\\ -0.079\end{subarray}$ $\bm{-0.17}\begin{subarray}{c}+0.079\\ -0.08\end{subarray}$ $\bm{-0.1}\begin{subarray}{c}+0.079\\ -0.08\end{subarray}$ $\bm{-0.11}\begin{subarray}{c}+0.078\\ -0.078\end{subarray}$ $\bm{-0.13}\begin{subarray}{c}+0.082\\ -0.081\end{subarray}$
Hadrons $+0.01\begin{subarray}{c}+0.1\\ -0.1\end{subarray}$ $-0.11\begin{subarray}{c}+0.17\\ -0.2\end{subarray}$ $-0.02\begin{subarray}{c}+0.1\\ -0.1\end{subarray}$ $+0.003\begin{subarray}{c}+0.02\\ -0.02\end{subarray}$ $-0.008\begin{subarray}{c}+0.02\\ -0.03\end{subarray}$ $+0.0006\begin{subarray}{c}+0.01\\ -0.01\end{subarray}$
String theory & supergravity $\bm{+0.3}\begin{subarray}{c}+0.14\\ -0.14\end{subarray}$ $\bm{+0.39}\begin{subarray}{c}+0.18\\ -0.18\end{subarray}$ $\bm{+0.15}\begin{subarray}{c}+0.13\\ -0.13\end{subarray}$ $\bm{+0.02}\begin{subarray}{c}+0.03\\ -0.02\end{subarray}$ $\bm{+0.04}\begin{subarray}{c}+0.04\\ -0.02\end{subarray}$ $+0.004\begin{subarray}{c}+0.01\\ -0.008\end{subarray}$
Perturbative methods $+0.1\begin{subarray}{c}+0.21\\ -0.17\end{subarray}$ $+0.03\begin{subarray}{c}+0.2\\ -0.2\end{subarray}$ $+0.09\begin{subarray}{c}+0.2\\ -0.2\end{subarray}$ $+0.01\begin{subarray}{c}+0.03\\ -0.02\end{subarray}$ $+0.01\begin{subarray}{c}+0.03\\ -0.03\end{subarray}$ $+0.008\begin{subarray}{c}+0.03\\ -0.01\end{subarray}$
Classical fields $-0.13\begin{subarray}{c}+0.29\\ -0.46\end{subarray}$ $-0.1\begin{subarray}{c}+0.34\\ -0.44\end{subarray}$ $-0.08\begin{subarray}{c}+0.2\\ -0.4\end{subarray}$ $-0.009\begin{subarray}{c}+0.03\\ -0.05\end{subarray}$ $-0.009\begin{subarray}{c}+0.04\\ -0.05\end{subarray}$ $-0.003\begin{subarray}{c}+0.02\\ -0.03\end{subarray}$
Collider physics $\bm{-0.17}\begin{subarray}{c}+0.15\\ -0.16\end{subarray}$ $\bm{-0.35}\begin{subarray}{c}+0.18\\ -0.19\end{subarray}$ $\bm{-0.21}\begin{subarray}{c}+0.15\\ -0.15\end{subarray}$ $-0.005\begin{subarray}{c}+0.02\\ -0.02\end{subarray}$ $\bm{-0.02}\begin{subarray}{c}+0.02\\ -0.03\end{subarray}$ $-0.004\begin{subarray}{c}+0.01\\ -0.02\end{subarray}$
Dynamics $-0.05\begin{subarray}{c}+0.3\\ -0.3\end{subarray}$ $-0.02\begin{subarray}{c}+0.3\\ -0.4\end{subarray}$ $-0.02\begin{subarray}{c}+0.25\\ -0.29\end{subarray}$ $+0.0006\begin{subarray}{c}+0.03\\ -0.03\end{subarray}$ $+0.003\begin{subarray}{c}+0.05\\ -0.05\end{subarray}$ $+0.001\begin{subarray}{c}+0.02\\ -0.02\end{subarray}$
Neutrinos & flavour physics $\bm{+0.24}\begin{subarray}{c}+0.17\\ -0.16\end{subarray}$ $+0.17\begin{subarray}{c}+0.19\\ -0.18\end{subarray}$ $\bm{+0.15}\begin{subarray}{c}+0.16\\ -0.15\end{subarray}$ $+0.02\begin{subarray}{c}+0.03\\ -0.02\end{subarray}$ $+0.02\begin{subarray}{c}+0.03\\ -0.02\end{subarray}$ $+0.006\begin{subarray}{c}+0.02\\ -0.009\end{subarray}$
Black holes $+0.05\begin{subarray}{c}+0.2\\ -0.1\end{subarray}$ $+0.18\begin{subarray}{c}+0.22\\ -0.2\end{subarray}$ $+0.02\begin{subarray}{c}+0.1\\ -0.1\end{subarray}$ $-0.007\begin{subarray}{c}+0.02\\ -0.02\end{subarray}$ $+0.008\begin{subarray}{c}+0.03\\ -0.02\end{subarray}$ $-0.005\begin{subarray}{c}+0.01\\ -0.02\end{subarray}$
Gauge theory & Grand Unification $-0.08\begin{subarray}{c}+0.2\\ -0.3\end{subarray}$ $-0.1\begin{subarray}{c}+0.3\\ -0.4\end{subarray}$ $-0.04\begin{subarray}{c}+0.2\\ -0.3\end{subarray}$ $-0.003\begin{subarray}{c}+0.03\\ -0.03\end{subarray}$ $-0.009\begin{subarray}{c}+0.04\\ -0.05\end{subarray}$ $-9e-05\begin{subarray}{c}+0.02\\ -0.02\end{subarray}$
Dark matter $-0.19\begin{subarray}{c}+0.2\\ -0.24\end{subarray}$ $\bm{-0.27}\begin{subarray}{c}+0.25\\ -0.27\end{subarray}$ $-0.16\begin{subarray}{c}+0.19\\ -0.24\end{subarray}$ $-0.01\begin{subarray}{c}+0.02\\ -0.04\end{subarray}$ $-0.02\begin{subarray}{c}+0.027\\ -0.043\end{subarray}$ $-0.007\begin{subarray}{c}+0.01\\ -0.03\end{subarray}$
Thermodynamics $+0.08\begin{subarray}{c}+0.3\\ -0.2\end{subarray}$ $+0.2\begin{subarray}{c}+0.4\\ -0.3\end{subarray}$ $+0.08\begin{subarray}{c}+0.3\\ -0.2\end{subarray}$ $+0.006\begin{subarray}{c}+0.04\\ -0.03\end{subarray}$ $+0.02\begin{subarray}{c}+0.06\\ -0.04\end{subarray}$ $+0.004\begin{subarray}{c}+0.03\\ -0.02\end{subarray}$
Cosmology $-0.01\begin{subarray}{c}+0.1\\ -0.2\end{subarray}$ $+0.03\begin{subarray}{c}+0.2\\ -0.2\end{subarray}$ $-0.03\begin{subarray}{c}+0.1\\ -0.2\end{subarray}$ $-0.009\begin{subarray}{c}+0.02\\ -0.03\end{subarray}$ $-0.004\begin{subarray}{c}+0.02\\ -0.03\end{subarray}$ $-0.006\begin{subarray}{c}+0.01\\ -0.02\end{subarray}$
Electroweak sector $-0.06\begin{subarray}{c}+0.1\\ -0.1\end{subarray}$ $-0.16\begin{subarray}{c}+0.17\\ -0.18\end{subarray}$ $-0.01\begin{subarray}{c}+0.1\\ -0.1\end{subarray}$ $+0.002\begin{subarray}{c}+0.02\\ -0.02\end{subarray}$ $-0.009\begin{subarray}{c}+0.02\\ -0.02\end{subarray}$ $+0.006\begin{subarray}{c}+0.02\\ -0.009\end{subarray}$
QCD $-0.04\begin{subarray}{c}+0.1\\ -0.2\end{subarray}$ $-0.12\begin{subarray}{c}+0.18\\ -0.2\end{subarray}$ $-0.04\begin{subarray}{c}+0.1\\ -0.1\end{subarray}$ $+0.0009\begin{subarray}{c}+0.02\\ -0.02\end{subarray}$ $-0.01\begin{subarray}{c}+0.022\\ -0.029\end{subarray}$ $+0.001\begin{subarray}{c}+0.02\\ -0.01\end{subarray}$
Quantum Field Theory $+0.04\begin{subarray}{c}+0.2\\ -0.2\end{subarray}$ $+0.12\begin{subarray}{c}+0.25\\ -0.21\end{subarray}$ $+0.04\begin{subarray}{c}+0.2\\ -0.2\end{subarray}$ $-0.006\begin{subarray}{c}+0.02\\ -0.03\end{subarray}$ $+0.0001\begin{subarray}{c}+0.03\\ -0.03\end{subarray}$ $-0.002\begin{subarray}{c}+0.01\\ -0.02\end{subarray}$
AdS/CFT $+0.005\begin{subarray}{c}+0.2\\ -0.2\end{subarray}$ $+0.08\begin{subarray}{c}+0.3\\ -0.2\end{subarray}$ $+0.009\begin{subarray}{c}+0.2\\ -0.2\end{subarray}$ $-0.007\begin{subarray}{c}+0.02\\ -0.03\end{subarray}$ $-0.001\begin{subarray}{c}+0.04\\ -0.03\end{subarray}$ $-0.003\begin{subarray}{c}+0.02\\ -0.02\end{subarray}$
Table 4: Effect of each variable on (a) the probability of having entered a new research area and (b) the probability of having exited a research area, for each model. The reference model uses entropy as the diversity measure $D$ and the magnitude of intellectual capital as a measure of power $P$. Values indicate the mean posterior effect size and the 95% credible interval. Significant effects are shown in bold.
Dep. variable Entered a new research area Exited a research area
Model Reference $D=\text{Stirling}$ $P=\text{Brokerage}$ Reference $D=\text{Stirling}$ $P=\text{Brokerage}$
Predictor
Intellectual capital (diversity) $\bm{+0.15}\begin{subarray}{c}+0.11\\ -0.11\end{subarray}$ $\bm{+0.14}\begin{subarray}{c}+0.099\\ -0.1\end{subarray}$ $\bm{+0.12}\begin{subarray}{c}+0.11\\ -0.11\end{subarray}$ $\bm{+1.1}\begin{subarray}{c}+0.14\\ -0.14\end{subarray}$ $\bm{+0.86}\begin{subarray}{c}+0.12\\ -0.12\end{subarray}$ $\bm{+0.96}\begin{subarray}{c}+0.14\\ -0.13\end{subarray}$
Social capital (diversity) $\bm{+0.2}\begin{subarray}{c}+0.1\\ -0.1\end{subarray}$ $\bm{+0.16}\begin{subarray}{c}+0.1\\ -0.1\end{subarray}$ $\bm{+0.18}\begin{subarray}{c}+0.1\\ -0.1\end{subarray}$ $+0.06\begin{subarray}{c}+0.1\\ -0.1\end{subarray}$ $+0.04\begin{subarray}{c}+0.1\\ -0.1\end{subarray}$ $+0.01\begin{subarray}{c}+0.1\\ -0.1\end{subarray}$
Social capital (power) $\bm{-0.12}\begin{subarray}{c}+0.1\\ -0.1\end{subarray}$ $\bm{-0.11}\begin{subarray}{c}+0.1\\ -0.1\end{subarray}$ $-0.07\begin{subarray}{c}+0.1\\ -0.1\end{subarray}$ $\bm{-0.31}\begin{subarray}{c}+0.11\\ -0.11\end{subarray}$ $\bm{-0.26}\begin{subarray}{c}+0.11\\ -0.11\end{subarray}$ $\bm{-0.18}\begin{subarray}{c}+0.1\\ -0.1\end{subarray}$
Stable affiliation $-0.14\begin{subarray}{c}+0.19\\ -0.19\end{subarray}$ $-0.14\begin{subarray}{c}+0.19\\ -0.19\end{subarray}$ $-0.15\begin{subarray}{c}+0.2\\ -0.19\end{subarray}$ $-0.11\begin{subarray}{c}+0.21\\ -0.21\end{subarray}$ $-0.07\begin{subarray}{c}+0.2\\ -0.2\end{subarray}$ $-0.15\begin{subarray}{c}+0.21\\ -0.21\end{subarray}$
Hadrons $-0.11\begin{subarray}{c}+0.31\\ -0.36\end{subarray}$ $-0.21\begin{subarray}{c}+0.33\\ -0.39\end{subarray}$ $-0.13\begin{subarray}{c}+0.3\\ -0.36\end{subarray}$ $-0.04\begin{subarray}{c}+0.3\\ -0.3\end{subarray}$ $-0.27\begin{subarray}{c}+0.39\\ -0.45\end{subarray}$ $-0.06\begin{subarray}{c}+0.3\\ -0.3\end{subarray}$
String theory & supergravity $\bm{+0.44}\begin{subarray}{c}+0.35\\ -0.32\end{subarray}$ $\bm{+0.53}\begin{subarray}{c}+0.36\\ -0.35\end{subarray}$ $\bm{+0.36}\begin{subarray}{c}+0.33\\ -0.31\end{subarray}$ $\bm{+0.46}\begin{subarray}{c}+0.36\\ -0.34\end{subarray}$ $\bm{+0.9}\begin{subarray}{c}+0.42\\ -0.42\end{subarray}$ $+0.23\begin{subarray}{c}+0.34\\ -0.25\end{subarray}$
Perturbative methods $+0.05\begin{subarray}{c}+0.4\\ -0.4\end{subarray}$ $+0.03\begin{subarray}{c}+0.4\\ -0.4\end{subarray}$ $+0.06\begin{subarray}{c}+0.4\\ -0.4\end{subarray}$ $-0.13\begin{subarray}{c}+0.33\\ -0.46\end{subarray}$ $-0.33\begin{subarray}{c}+0.47\\ -0.57\end{subarray}$ $-0.1\begin{subarray}{c}+0.3\\ -0.4\end{subarray}$
Classical fields $+0.26\begin{subarray}{c}+1.2\\ -0.63\end{subarray}$ $+0.4\begin{subarray}{c}+1.5\\ -0.75\end{subarray}$ $+0.26\begin{subarray}{c}+1.2\\ -0.63\end{subarray}$ $-0.43\begin{subarray}{c}+0.64\\ -1.6\end{subarray}$ $-0.51\begin{subarray}{c}+0.85\\ -1.3\end{subarray}$ $-0.27\begin{subarray}{c}+0.48\\ -1.2\end{subarray}$
Collider physics $\bm{-0.35}\begin{subarray}{c}+0.33\\ -0.35\end{subarray}$ $\bm{-0.52}\begin{subarray}{c}+0.34\\ -0.35\end{subarray}$ $\bm{-0.37}\begin{subarray}{c}+0.33\\ -0.35\end{subarray}$ $+0.06\begin{subarray}{c}+0.31\\ -0.25\end{subarray}$ $-0.2\begin{subarray}{c}+0.35\\ -0.41\end{subarray}$ $+0.02\begin{subarray}{c}+0.26\\ -0.23\end{subarray}$
Dynamics $-0.18\begin{subarray}{c}+0.6\\ -0.85\end{subarray}$ $-0.18\begin{subarray}{c}+0.66\\ -0.89\end{subarray}$ $-0.16\begin{subarray}{c}+0.58\\ -0.83\end{subarray}$ $-0.02\begin{subarray}{c}+0.6\\ -0.7\end{subarray}$ $+0.14\begin{subarray}{c}+1.1\\ -0.86\end{subarray}$ $-0.006\begin{subarray}{c}+0.6\\ -0.6\end{subarray}$
Neutrinos & flavour physics $+0.14\begin{subarray}{c}+0.37\\ -0.29\end{subarray}$ $+0.1\begin{subarray}{c}+0.4\\ -0.3\end{subarray}$ $+0.11\begin{subarray}{c}+0.35\\ -0.29\end{subarray}$ $-0.19\begin{subarray}{c}+0.27\\ -0.35\end{subarray}$ $-0.33\begin{subarray}{c}+0.38\\ -0.44\end{subarray}$ $-0.25\begin{subarray}{c}+0.28\\ -0.37\end{subarray}$
Black holes $-0.07\begin{subarray}{c}+0.3\\ -0.4\end{subarray}$ $+0.006\begin{subarray}{c}+0.4\\ -0.4\end{subarray}$ $-0.08\begin{subarray}{c}+0.3\\ -0.4\end{subarray}$ $+0.16\begin{subarray}{c}+0.48\\ -0.32\end{subarray}$ $\bm{+0.78}\begin{subarray}{c}+0.59\\ -0.56\end{subarray}$ $+0.12\begin{subarray}{c}+0.44\\ -0.28\end{subarray}$
Gauge theory & Grand Unification $+0.003\begin{subarray}{c}+0.6\\ -0.6\end{subarray}$ $+0.009\begin{subarray}{c}+0.7\\ -0.6\end{subarray}$ $+0.02\begin{subarray}{c}+0.6\\ -0.6\end{subarray}$ $-0.11\begin{subarray}{c}+0.52\\ -0.78\end{subarray}$ $-0.12\begin{subarray}{c}+0.8\\ -0.88\end{subarray}$ $-0.06\begin{subarray}{c}+0.4\\ -0.6\end{subarray}$
Dark matter $\bm{-0.65}\begin{subarray}{c}+0.56\\ -0.55\end{subarray}$ $\bm{-0.71}\begin{subarray}{c}+0.56\\ -0.57\end{subarray}$ $\bm{-0.64}\begin{subarray}{c}+0.55\\ -0.56\end{subarray}$ $+0.02\begin{subarray}{c}+0.4\\ -0.4\end{subarray}$ $-0.002\begin{subarray}{c}+0.5\\ -0.5\end{subarray}$ $+0.02\begin{subarray}{c}+0.4\\ -0.3\end{subarray}$
Thermodynamics $-0.07\begin{subarray}{c}+0.6\\ -0.7\end{subarray}$ $-0.04\begin{subarray}{c}+0.6\\ -0.6\end{subarray}$ $-0.05\begin{subarray}{c}+0.6\\ -0.6\end{subarray}$ $+0.05\begin{subarray}{c}+0.7\\ -0.6\end{subarray}$ $+0.35\begin{subarray}{c}+1.2\\ -0.77\end{subarray}$ $+0.04\begin{subarray}{c}+0.6\\ -0.5\end{subarray}$
Cosmology $+0.004\begin{subarray}{c}+0.4\\ -0.3\end{subarray}$ $+0.06\begin{subarray}{c}+0.4\\ -0.4\end{subarray}$ $-0.004\begin{subarray}{c}+0.3\\ -0.3\end{subarray}$ $+0.15\begin{subarray}{c}+0.5\\ -0.33\end{subarray}$ $\bm{+0.56}\begin{subarray}{c}+0.6\\ -0.54\end{subarray}$ $+0.1\begin{subarray}{c}+0.44\\ -0.29\end{subarray}$
Electroweak sector $+0.1\begin{subarray}{c}+0.33\\ -0.28\end{subarray}$ $+0.06\begin{subarray}{c}+0.3\\ -0.3\end{subarray}$ $+0.12\begin{subarray}{c}+0.35\\ -0.29\end{subarray}$ $+0.05\begin{subarray}{c}+0.3\\ -0.3\end{subarray}$ $-0.03\begin{subarray}{c}+0.4\\ -0.4\end{subarray}$ $+0.1\begin{subarray}{c}+0.35\\ -0.24\end{subarray}$
QCD $-0.07\begin{subarray}{c}+0.31\\ -0.35\end{subarray}$ $-0.12\begin{subarray}{c}+0.33\\ -0.38\end{subarray}$ $-0.07\begin{subarray}{c}+0.3\\ -0.4\end{subarray}$ $+0.02\begin{subarray}{c}+0.3\\ -0.3\end{subarray}$ $+0.008\begin{subarray}{c}+0.4\\ -0.4\end{subarray}$ $+0.03\begin{subarray}{c}+0.3\\ -0.3\end{subarray}$
Quantum Field Theory $+0.3\begin{subarray}{c}+0.61\\ -0.42\end{subarray}$ $+0.41\begin{subarray}{c}+0.65\\ -0.49\end{subarray}$ $+0.31\begin{subarray}{c}+0.62\\ -0.43\end{subarray}$ $-0.03\begin{subarray}{c}+0.4\\ -0.4\end{subarray}$ $+0.22\begin{subarray}{c}+0.61\\ -0.48\end{subarray}$ $-0.01\begin{subarray}{c}+0.4\\ -0.4\end{subarray}$
AdS/CFT $+0.17\begin{subarray}{c}+0.64\\ -0.45\end{subarray}$ $+0.26\begin{subarray}{c}+0.72\\ -0.5\end{subarray}$ $+0.18\begin{subarray}{c}+0.65\\ -0.45\end{subarray}$ $-0.05\begin{subarray}{c}+0.4\\ -0.5\end{subarray}$ $+0.21\begin{subarray}{c}+0.74\\ -0.57\end{subarray}$ $-0.03\begin{subarray}{c}+0.4\\ -0.5\end{subarray}$
A.8 Exploring individual careers
Tables 5 and 6 show the 20 most conservative and least conservative physicists. Measures of the diversity of intellectual capital ($D(\bm{I_{a}})$) and social capital ($D(\bm{S_{a}})$) are included. Table 5 shows the physicists who have been the most conservative towards their research agenda. Interestingly, 15 out of 20 are specialized in “Collider physics”, in line with the above observation that this particular research area is more specialized.
Table 6 shows the 10 physicists with the highest change scores. Looking into the share of attention dedicated by each scientist to each area, one can see that high change scores are usually reached, unsurprisingly, when physicists exit a research area and enter a new one in which they had almost no prior interest.
In order to make sure that these cases of high migration were not false positives due to the failure to properly disambiguate authors, we have inspected each of them manually, by comparing physicists’ Inspire HEP profiles with data from their personal websites, as well as their university or LinkedIn profiles. We found one of them to be a false positive due to failing author disambiguation, namely “T. Fuchs”, i.e. the author with the highest change score. In the case of S. D. Katz, certain publications were excluded due to inaccurate classifications in the database, which has led to overestimate their initial distribution of dark matter research. All other cases are accurate. For instance, on G. Ferretti’s university webpage, one can read: “Gabriele Ferretti joined Chalmers in 1999. For the first ten years he has been mainly involved in research in quantum field theory and string theory with emphasis on their formal aspects. Since the start of the LHC he has also become interested in phenomenological applications of quantum field theory and model building”  [^20] https://www.chalmers.se/en/persons/ferretti/, accessed on 11/10/2023.
Comparing tables 5 and 6, one can see that physicists with high change scores tend to have much more diverse intellectual and social capital. In other words, such physicists have had more opportunities to adapt or revise their research agenda. This is in line with the observation that change is enhanced by the diversity of cognitive resources one has access to.
Table 5: Physicists with the lowest change scores $c_{a}$. $D(\bm{I_{a}})$ and $D(\bm{S_{a}})$ measure the diversity of intellectual and social capital. Numbers in parentheses indicate the share of attention dedicated to each research area. Asterisks ($\ast$) indicate physicists with a permanent position.
Physicist
$c_{a}$ $D(\bm{I_{a}})$ $D(\bm{S_{a}})$
Previous main area
Current main area
J. Huston ($\ast$)
0.04 2.96 5.50
Collider physics (0.61)
Collider physics (0.60)
U. D. Alesio ($\ast$)
0.05 1.68 4.50
Collider physics (0.87)
Collider physics (0.88)
S. Schumann
0.06 4.15 7.90
Collider physics (0.46)
Collider physics (0.45)
A. V. Lipatov
0.06 2.88 4.28
Collider physics (0.68)
Collider physics (0.68)
Andreas Metz
0.06 2.01 3.61
Collider physics (0.85)
Collider physics (0.89)
I. Vitev
0.07 2.95 5.49
Collider physics (0.72)
Collider physics (0.73)
Xin Nian Wang ($\ast$)
0.07 2.61 4.06
Collider physics (0.77)
Collider physics (0.75)
S. Goswami
0.07 2.15 5.22
Neutrinos & flavour physics (0.83)
Neutrinos & flavour physics (0.79)
Y. Mehtar Tani
0.07 2.75 5.17
Collider physics (0.75)
Collider physics (0.75)
M. Trigiante
0.07 5.11 6.42
String theory & supergravity (0.55)
String theory & supergravity (0.56)
F. Murgia ($\ast$)
0.07 1.84 4.57
Collider physics (0.86)
Collider physics (0.87)
B. Kayser ($\ast$)
0.07 2.07 4.94
Neutrinos & flavour physics (0.76)
Neutrinos & flavour physics (0.77)
J. F. Owens ($\ast$)
0.07 2.01 5.42
Collider physics (0.78)
Collider physics (0.80)
A. Bacchetta
0.07 2.07 3.19
Collider physics (0.84)
Collider physics (0.81)
P. M. Nadolsky ($\ast$)
0.07 3.19 5.28
Collider physics (0.62)
Collider physics (0.59)
D. Martelli
0.07 4.42 5.97
String theory & supergravity (0.57)
String theory & supergravity (0.58)
C. Pisano
0.08 1.77 4.53
Collider physics (0.87)
Collider physics (0.88)
M. Strikman ($\ast$)
0.08 2.77 4.99
Collider physics (0.77)
Collider physics (0.75)
J. Nemchik ($\ast$)
0.08 2.67 4.69
Collider physics (0.80)
Collider physics (0.78)
Zhi Zhong Xing ($\ast$)
0.08 3.44 8.17
Neutrinos & flavour physics (0.71)
Neutrinos & flavour physics (0.65)
Table 6: Physicists with the highest change scores $c_{a}$. $D(\bm{I_{a}})$ and $D(\bm{S_{a}})$ measure the diversity of intellectual and social capital. Numbers in parentheses indicate the share of attention dedicated to each research area during each time-period. Asterisks ($\ast$) indicate physicists with a permanent position.
Physicist
$c_{a}$ $D(\bm{I_{a}})$ $D(\bm{S_{a}})$
Previous main area
Current main area
T. Fuchs
0.93 4.38 6.50
QCD (0.53$\to$0.00)
Neutrinos & flavour physics (0.00$\to$0.64)
N. G. Sanchez ($\ast$)
0.88 7.52 13.08
Black holes (0.37$\to$0.03)
Dark matter (0.00$\to$0.57)
V. V. Braguta ($\ast$)
0.88 5.03 8.04
Hadrons (0.45$\to$0.00)
QCD (0.05$\to$0.60)
G. B. Cleaver ($\ast$)
0.83 6.34 12.22
String theory & supergravity (0.40$\to$0.00)
Cosmology (0.03$\to$0.62)
J. Martin Camalich
0.82 2.34 4.57
QCD (0.78$\to$0.03)
Hadrons (0.06$\to$0.38)
Kwei Chou Yang ($\ast$)
0.81 3.58 7.80
Hadrons (0.63$\to$0.03)
Dark matter (0.00$\to$0.78)
S. D. Katz ($\ast$)
0.80 6.50 10.85
Dark matter (0.42$\to$0.03)
QCD (0.06$\to$0.47)
G. Ferretti ($\ast$)
0.78 5.22 6.18
String theory & supergravity (0.52$\to$0.00)
Electroweak sector (0.01$\to$0.58)
K. Bhattacharya
0.78 10.79 10.59
Neutrinos & flavour physics (0.20$\to$0.00)
Cosmology (0.09$\to$0.37)
H. Ebrahim ($\ast$)
0.78 3.90 6.81
Black holes (0.44$\to$0.06)
AdS/CFT (0.02$\to$0.34)
K. Narayan
0.77 3.83 7.13
String theory & supergravity (0.62$\to$0.00)
AdS/CFT (0.03$\to$0.52)
Mario Campanelli
0.77 2.04 4.48
Neutrinos & flavour physics (0.79$\to$0.06)
Electroweak sector (0.09$\to$0.64)
V. K. Oikonomou ($\ast$)
0.76 8.05 6.18
Thermodynamics (0.24$\to$0.02)
Cosmology (0.00$\to$0.47)
K. Sasaki
0.76 3.14 8.87
Collider physics (0.68$\to$0.06)
Thermodynamics (0.00$\to$0.27)
H. L. Verlinde ($\ast$)
0.74 5.32 8.88
String theory & supergravity (0.60$\to$0.04)
AdS/CFT (0.02$\to$0.35)
Bin Wu
0.74 5.02 8.01
QCD (0.41$\to$0.02)
Collider physics (0.15$\to$0.56)
Eiji Nakano ($\ast$)
0.74 6.90 10.13
QCD (0.39$\to$0.06)
Hadrons (0.04$\to$0.78)
A. T. Suzuki ($\ast$)
0.73 4.36 6.43
Perturbative methods (0.61$\to$0.04)
Neutrinos & flavour physics (0.01$\to$0.39)
A. Mariotti
0.73 3.81 4.91
String theory & supergravity (0.63$\to$0.01)
Electroweak sector (0.01$\to$0.38)
I. Scimemi
0.73 6.77 8.60
Hadrons (0.36$\to$0.00)
Collider physics (0.09$\to$0.81)
A.9 Replication corpus
For purposes of replication, certain analyses have been reproduced on the ACL anthology corpus of Computational Linguistics research [15]. A comparison with the HEP corpus in shown in Table 7. The ACL corpus is smaller, presumably due to a lack of exhaustivity. The requirement on the quantity of publications has been lowered (three per time period instead of five) in order to maintain a reasonable size of cohort. The choice of time periods is arbitrary (i.e. it was not informed by the context). Only 14 topics were deemed to carry scientific content valid upon manual inspection.
Table 7: Comparison of the high-energy physics corpus and the computational linguistics corpus.
Cohort High-energy physics Computational Linguistics
Cohort size ($N$) 2 19521952\,1952 195 $578$
Number of abstracts ($D$) 186 162186162186\,162186 162 58 2895828958\,28958 289
Vocabulary size ($V$) 4 75147514\,7514 751 5 96659665\,9665 966
Valid topics ($K$) 16 14
Initial time period 2000–2009 2002–2011
Late time period 2015–2019 2012–2022
Author disambiguation Yes No
Minimum publications $5$ $3$
The transfers of attention are shown in Figure 16. Compared to the high-energy physics corpus, it features significant disruptions (e.g., the emergence of new topics, such as “deep learning”, “sentiment analysis” and “embeddings & pre-trained models”).
Figure 16: Aggregate transfers of attention across research areas in the ACL anthology corpus of Computational Linguistics, between 2002-2011 (to the left) and 2012-2022 (to the right). Widths of flows are proportional to $\sum_{a}X_{ak}\theta_{akk^{\prime}}$. Insignificant transfers (that happen less than expected by chance alone assuming uniform mixing) are transparent.